MULTIPHOTON IONIZATION AND DISSOCIATION OF DIAZIRINE

\underline{ANDREW K. MOLLNER}, I. FEDOROV, L. KOZIOL, A. I. KRYLOV, H. REISLER, Department of Chemistry, University of Southern California, Los Angeles, CA 90089.

Multiphoton ionization and dissociation processes in diazirine have been studied experimentally via 304-325 nm two-photon absorption, and theoretically by using the EOM-CCSD and B3LYP methods. The electronic structure calculations indicate the strongest one-photon absorption is to the 2^{1}A_{1}(3p_{x}←n) Rydberg state. However, in two-photon absorption at comparable energies the first photon excites the low-lying 1^{1}B_{2}(Π^{∗}←n) valence state, from which the strongest absorption is to the dissociative valence 1^{1}A_{2}(Π^{∗}←σ_{NN}) state. In the experimental studies, resonance enhanced multiphoton ionization (REMPI) experiments show no ions at the parent diazirine mass but only CH_{2}^{+} ions from dissociative photoionization. It is proposed that weak one-photon absorption to the 1^{1}B_{2} state is immediately followed by more efficient absorption of another photon to reach the 1^{1}A_{2} state from which competition between ionization and fast dissociation takes place. Strong signals of CH^{+} ions are also detected and assigned to 2+1 of CH fragments. Velocity map CH^{+} images show that CH fragments are born with substantial translational energy indicating that they arise from absorption of two photons in diazirine. It is argued that two-photon processes via the 1^{1}B_{2} intermediate state are very efficient in this wavelength range, leading predominantly to dissociation of diazirine from the 1^{1}A_{2} state. The most likely route to CH(X) formation is isomerization to isodiazirine followed by dissociation to CH + HN_{2}.